KLEIN BOTTLES OF GENUS 4

by Adrian Jones A Senior Honors Thesis Submitted to the Faculty of the University of Utah in Partial Fulfillment of the Requirements for the Honors Degree of Bachelor of Arts in Mathematics

APPROVED:

Aaron Bertram  Graeme Milton
Supervisor  Chair, Mathematics
Fletcher Gross  Martha S. Bradley
Departmental Honors Advisor  Director, Honors Program

August 20, 2003

ABSTRACT

Blah blah blah blah blah blah blah blah.

Blah blah blah blah blah blah blah blah.

TABLE OF CONTENTS

Abstract ii 1 Introduction 1 2 Klein Bottles of Genus 2 2 3 Klein Bottles of Genus 3 3 4 Klein Bottles of Genus 4 4

Chapter 1 Introduction

The standard treatise in this field is Menapace’s small monograph [1]. Blah blah blah blah blah blah blah blah.

Chapter 2 Klein Bottles of Genus 2

Blah blah blah blah blah blah blah blah.

Bibliography

[1] John Menapace. Letter in a Klein bottle: photographs. Jargon Society, East Haven, CT, USA, 1984. ISBN 0-912330-56-2. 8 + 60 + 16 pp. LCCN TR647 .M45 1984. Introduction by Donald B. Kuspit. Afterword by Jonathan Williams.

Name of Candidate:  Adrian Jones
Birth Date:  24 July 1847
Birth Place:  Nauvoo, IL
Address:  1234 South Main, Salt Lake City, UT